You are valuing a block of three units with the following Capital Value (CIV) and Estimated Annual Value (EAV) provided in the below. Adopting the relevant provision in Value for a larger Improved table the Act for property calculate the individual Site Values for each unit Site an assessed total Site Value of $1,800,000 (round to the nearest dollar and provide the answers in your booklet) Units Capital Improved Value $600,000 $900,000 $1,200,000 Site Value Estimated Annual Value Unit 1 Unit 2 Unit 3 $45,000 $54.000 Total $129,000 $1,800.000 (3 marks) 8. What is a supplementary valuation made under the Act? Briefly describe in your own words, and provide two circumstances requiring the making of (3 marks) a supplementary valuation. 9. Briefly describe three functions of the Valuer-General under the Act (3 marks) 10. State if full what the acronym AVPCC stands for and explain how or on 2 marks) what basis is the AVPCC determined or allocated? 11.According to the Australian Tax Office (ATO) terminology, provide an example of a property that would be Goods and Services Tax (GST) free. 1 mark) 12. What is the valuation base used to apply the Fire Services Property Levy? (1 mark) 13. What is the valuation base used to assess Land Tax in Victoria? (1 mark)
Answer #1

Answer 7 -

The individual site value for each unit can be calculated by taking into account 'Estimated Annual Value' of each unit.

Statement showing calculation of Individual Site value for each unit:

Unit Estimated Annual Value ($) Site Value ($) Working/Calculations
1 30,000 418,605 =1,800,000*30,000/129,000
2 45,000 627,907 =1,800,000*45,000/129,000
3 54,000 753,488 =1,800,000*54,000/129,000
129,000